# Low-Rank Quantization-Aware Training for LLMs

### v2

@@ Appendix A Extended results
In this section, we provide extended results and present some additional ablation studies.
Table A1: LM-eval weight-only quantization results for LLaMA-2/3 and Mistral models. We report zero-shot accuracy of 6 tasks (higher is better).
Model # Bits Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
LLaMA-2 7B FP16 77.74 79.11 69.14 74.58 46.25 75.98 70.47
@@
LR-QAT (ours) 73.30 78.07 67.72 71.46 43.77 73.53 67.98
LLaMA-2 13B FP16 80.55 80.52 72.22 77.44 48.98 79.38 73.18
@@
LR-QAT (ours) 79.91 80.25 71.03 77.48 50.68 79.45 73.13
@@
LR-QAT (ours) 79.69 78.78 68.98 73.99 44.88 76.89 70.54
LLaMA-3 8B FP16 81.44 80.79 72.85 77.74 53.33 79.16 74.22
@@
LR-QAT (ours) 80.09 80.41 73.88 76.89 50.51 78.30 73.35
@@
Mistral 7B FP16 83.58 82.10 73.88 79.59 53.92 81.07 75.69
@@
LR-QAT (ours) 80.12 81.83 73.40 76.73 49.66 80.25 73.67
@@
LR-QAT (ours) 81.38 80.41 70.32 73.15 45.39 78.00 71.44
Table A2: A comparison between min-max and the best range setting used for round-to-nearest (RTN) initialization for LLaMA-2/3 and Mistral models. We report WikiText-2 test set perplexity (lower is better) and average zero-shot accuracy (higher is better).
Model # Bits RTN init. WikiText-2 $\downarrow$ Zero-shot accuracy $\uparrow$
LLaMA-2 7B FP16 5.47 70.47
W4pc min-max 7.14 66.41
best ($L^{3.5}$) 6.14 68.88
W3pc min-max 1.9e4 35.71
best ($L^{3.5}$) 26.73 43.87
W4g128 best = min-max 5.78 69.75
W3g128 min-max 8.22 64.07
best ($L^{4}$) 7.61 63.20
LLaMA-2 13B FP16 4.88 73.18
W4pc min-max 5.40 72.19
best ($L^{3.5}$) 5.21 71.73
W3pc min-max 2.3e3 37.85
best ($L^{5}$) 8.71 55.01
W4g128 best = min-max 5.04 72.94
W3g128 min-max 6.14 66.81
best ($L^{5}$) 6.20 67.60
LLaMA-3 8B FP16 6.14 74.22
W4pc min-max 10.53 67.44
best ($L^{3.5}$) 7.53 72.19
W3pc min-max 1.6e5 35.78
best ($L^{3.5}$) 34.10 47.46
W4g128 min-max 6.99 72.95
best ($L^{4}$) 6.96 72.30
W3g128 min-max 29.38 54.54
best ($L^{5}$) 15.11 57.74
Mistral 7B FP16 5.25 75.69
W4pc min-max 6.33 71.84
best ($L^{4}$) 5.91 73.44
W3pc min-max 3.2e3 36.78
best ($L^{4}$) 9.49 64.58
W4g128 min-max 5.51 74.98
best ($L^{5}$) 5.49 75.07
W3g128 min-max 7.22 68.35
best ($L^{5}$) 6.77 69.35
@@ Appendix B Experimental details
Learning rate for $\bm{A}$, $\bm{B}$$\{10^{-5},10^{-4},10^{-3},10^{-2}\}$
Learning rate for quantization scale ($\mathbf{s}$) $\{0{{\textsuperscript{*}}},10^{-5}\}$
@@
Learning rate schedule for quantization scale ($\mathbf{s}$) linear (with warmup)
@@
Weight decay for quantization scale ($\mathbf{s}$) 00
Adam $\beta_{1}$ $0.9$
Adam $\beta_{2}$ $0.95$
@@
Table B1: Common hyperparameters used for experiments. *Is equivalent to freezing the quantization scale obtained after initial range estimation ($\mathbf{s}=\mathbf{s_{0}}$).
@@
We experiment with both weight-only and weight-activation quantization. The default settings are INT4 / INT3 per-channel (denoted ‘pc’) and group-wise weight quantization with a group size of 128 (denoted ‘g128’). We always use symmetric uniform affine quantization. We quantize all linear layers, except the classification head. LayerNorm and Embedding layers are always kept at full precision. In weight-activation quantization, defaults are INT4 per-channel weight and per-token activation quantization [14]. Following OmniQuant [55] we quantize all inputs to matmuls with exception of the softmax output and additionally quantize the KV-cache as in LLM-QAT [44].
@@
We implement our method in PyTorch [53] and use training and evaluation pipelines from HuggingFace libraries [20, 38, 64]. For zero-shot evaluation, we use the LM Evaluation Harness framework [19].
@@
We apply our methods to all linear layers in the attention blocks (both in self-attention and in the feed-forward network). We only train low-rank auxiliary matrices $\bm{A}$, $\bm{B}$ and the quantization parameters $\mathbf{s}$. Specifically, we freeze the token embeddings, the final classification heads and LayerNorm parameters.
We train on a small subset of SlimPajama [57], which is a close open-source replica of the dataset used for pre-training LLaMA models. We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set [^4] Specifically, we use the English subset of Wiki40b, https://huggingface.co/datasets/wiki40b, that contains cleaned-up text of English Wikipedia and training/validation splits. (512 sequences). For weight-only and weight-activation quantization results, we train for $10^{4}$ steps (but for LLaMA-3 and Mistral we only train for $5000$ steps due to constrained resources). For ablation studies, we always use shorter training of $10^{3}$ steps. The rest of the hyperparameters and their search spaces are listed in Table B1
@@
Each experiment for which we reported results, was executed on a single NVidia A100 GPU equipped with 80GB of VRAM. LLaMA-2 7B and 13B experiments needed respectively around 3 and 5 days for 10k training steps experiments. Mistral 7B and LLaMA-3 8B needed around 1.6 days for 5k training steps experiments. For obtaining all the results in the paper, including the ablations, we needed 107 GPU days (A100). Including preliminary experiments that did not make it in the final paper and hyperparameter turning we estimate the total compute costs of this research to approximately 500 GPU days.

### v3

@@ Appendix A Extended results
In this section, we provide additional detailed results.
Table A1: LM-eval weight-only quantization results for LLaMA-1 7B. We report zero-shot accuracy of 6 tasks (higher is better). §Uses asymmetric weight quantization.
# Bits Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
FP16 75.05 79.16 70.01 72.85 44.80 76.21 69.68
W4 pc RTN 73.18 78.78 69.14 71.38 44.37 74.22 68.51
GPTQ§ 67.70 76.00 66.70 66.90 43.00 69.40 64.95
LSQ (our impl.) 73.91 78.24 69.22 70.84 43.26 74.74 68.37
PEQA (our impl.) 74.71 78.29 70.09 70.33 42.24 75.27 68.49
LR-QAT (ours) 74.13 78.29 70.01 71.21 42.41 75.16 68.54
W4 g128 RTN 74.77 78.51 70.64 71.30 43.60 74.74 68.93
LSQ (our impl.) 75.90 79.22 70.01 71.42 43.34 75.15 69.17
PEQA (our impl.) 75.75 79.17 70.17 70.75 43.60 75.71 69.19
LR-QAT (ours) 75.29 78.62 69.61 71.59 44.11 75.67 69.15
W3 pc RTN 58.93 70.40 55.72 55.01 32.17 55.75 54.66
LSQ (our impl.) 71.35 77.97 68.82 66.33 40.10 73.14 66.29
PEQA (our impl.) 72.69 77.15 65.90 68.27 38.91 71.60 65.75
LR-QAT (ours) 73.24 78.18 67.40 67.47 40.53 72.77 66.60
W3 g128 RTN 69.48 76.33 64.40 64.44 38.65 67.67 63.50
LSQ (our impl.) 71.04 77.97 68.11 68.27 40.44 73.37 66.53
PEQA (our impl.) 71.65 78.24 68.51 68.18 40.10 73.30 66.66
LR-QAT (ours) 72.84 78.02 67.40 68.52 41.04 73.04 66.81
W2 pc RTN§ 43.24 52.61 51.62 27.48 26.28 26.26 37.92
PEQA (our impl.)§ 67.19 74.32 61.80 59.97 32.42 63.28 59.83
LR-QAT (ours)§ 68.07 74.27 64.80 61.03 36.95 65.48 61.77
W2 g128 RTN§ 41.16 58.05 50.20 34.97 23.12 30.93 39.74
PEQA (our impl.)§ 67.40 74.21 62.04 60.98 35.32 64.16 60.69
LR-QAT (ours)§ 67.65 74.86 61.80 62.37 37.03 65.90 61.60
Table A2: LM-eval weight-only quantization results for LLaMA-2 7B. We report zero-shot accuracy of 6 tasks (higher is better). §Uses asymmetric weight quantization.
# Bits Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
FP16 77.74 79.11 69.14 74.58 46.25 75.98 70.47
@@
LR-QAT (ours) 76.61 77.31 68.98 72.05 42.58 74.20 68.62
W2 pc RTN§ 39.57 50.71 50.20 26.73 26.28 25.63 36.52
PEQA (our impl.)§ 64.98 71.33 58.09 55.22 33.11 61.66 57.40
LR-QAT (ours)§ 69.88 73.07 62.90 56.57 34.04 63.71 60.03
W2 g128 RTN§ 39.08 55.93 50.99 29.84 24.06 27.71 37.94
PEQA (our impl.)§ 69.51 73.83 63.46 58.21 34.90 64.53 60.74
LR-QAT (ours)§ 70.18 73.83 63.69 60.27 34.98 67.24 61.70
Table A3: LM-eval weight-only quantization results for LLaMA-2 13B. We report zero-shot accuracy of 6 tasks (higher is better). §Uses asymmetric weight quantization. ✕ denotes out of memory.
# Bits Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
FP16 80.55 80.52 72.22 77.44 48.98 79.38 73.18
@@
LR-QAT (ours) 80.15 80.09 72.06 77.65 49.91 79.28 73.19
W4 g128 RTN 81.10 79.82 72.38 76.73 49.06 78.52 72.94
LSQ (our impl.)
PEQA (our impl.) 80.28 80.63 71.74 76.14 48.38 79.62 72.80
LR-QAT (ours) 80.73 80.30 71.74 76.14 49.06 79.51 72.91
@@
LR-QAT (ours) 79.79 79.60 70.64 74.24 46.76 78.00 71.51
W2 pc RTN§ 38.35 48.97 48.54 27.78 27.99 25.97 36.27
PEQA (our impl.)§ 71.19 74.48 59.67 62.96 37.29 68.14 62.29
LR-QAT (ours)§ 72.60 76.61 66.22 66.67 39.76 69.79 65.28
W2 g128 RTN§ 50.12 57.29 50.36 34.43 22.18 32.35 41.12
PEQA (our impl.)§ 70.34 76.88 66.85 64.94 38.74 70.66 64.74
LR-QAT (ours)§ 75.47 77.86 65.98 67.63 41.04 72.50 66.75
Table A4: LM-eval weight-only quantization results for LLaMA-3 8B. We report zero-shot accuracy of 6 tasks (higher is better). §Uses asymmetric weight quantization.
# Bits Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
FP16 81.44 80.79 72.85 77.74 53.33 79.16 74.22
@@
LR-QAT (ours) 81.62 79.98 72.85 78.32 52.05 78.19 73.84
W4 g128 RTN 79.48 79.27 73.56 75.08 48.81 77.61 72.30
LSQ (our impl.) 80.24 80.36 73.01 77.23 50.43 78.60 73.31
PEQA (our impl.) 80.98 80.14 72.61 76.18 49.57 78.45 72.99
LR-QAT (ours) 80.40 80.90 73.48 77.44 51.11 78.60 73.66
@@
W2 pc RTN§ 44.89 48.97 47.51 25.29 27.73 26.41 36.80
PEQA (our impl.)§ 62.54 72.96 59.51 56.82 33.70 60.77 57.72
LR-QAT (ours)§ 65.81 72.31 64.72 53.54 33.45 61.11 58.49
W2 g128 RTN§ 38.47 53.32 51.78 28.75 22.70 26.81 36.97
PEQA (our impl.)§ 58.53 72.52 61.96 57.41 35.32 61.14 57.81
LR-QAT (ours)§ 67.98 73.83 62.83 56.94 36.43 64.77 60.46
Table A5: LM-eval weight-only quantization results for Mistral 7B. We report zero-shot accuracy of 6 tasks (higher is better). §Uses asymmetric weight quantization.
# Bits Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
FP16 83.58 82.10 73.88 79.59 53.92 81.07 75.69
@@
LR-QAT (ours) 81.99 81.28 73.56 78.20 51.02 80.57 74.44
W4 g128 RTN 84.16 81.77 74.43 77.95 51.71 80.42 75.07
LSQ (our impl.) 80.40 80.63 73.16 76.26 48.72 78.24 72.90
PEQA (our impl.) 80.89 81.72 73.80 75.42 48.46 79.76 73.34
LR-QAT (ours) 83.55 81.61 74.51 78.28 52.90 80.84 75.28
@@
LR-QAT (ours) 81.71 80.90 70.48 75.08 47.78 78.50 72.41
W2 pc RTN§ 39.42 51.03 49.49 26.22 26.71 26.67 36.59
PEQA (our impl.)§ 74.34 75.95 66.22 62.21 38.48 69.68 64.48
LR-QAT (ours)§ 75.08 76.39 65.98 64.48 39.59 69.14 65.11
W2 g128 RTN§ 54.98 56.91 51.70 31.27 23.21 29.70 41.30
PEQA (our impl.)§ 75.41 75.84 64.96 65.45 38.23 68.00 64.65
LR-QAT (ours)§ 75.81 77.15 64.64 63.59 39.51 70.89 65.27
Table A6: LM-eval weight and activation quantization results for LLaMA models. We report zero-shot accuracy of 6 tasks (higher is better). §Uses asymmetric weight quantization.
Model # Bits (W-A-KV) Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
LLaMA-1 7B FP16 75.05 79.16 70.01 72.85 44.80 76.21 69.68
4-8-8 RTN 71.35 76.66 66.46 66.84 41.55 72.10 65.83
SmoothQuant 71.00 76.00 66.00 67.40 42.80 67.80 65.17
LLM-QAT 74.60 77.50 67.70 70.20 45.60 73.50 68.18
PEQA (our impl.) 74.86 78.24 70.01 70.12 42.83 75.14 68.53
LR-QAT (ours) 73.76 78.51 71.19 71.09 41.81 75.10 68.58
4-8-4 RTN 68.81 75.46 62.12 62.46 39.51 68.33 62.78
SmoothQuant 54.70 55.40 51.50 43.90 27.70 38.90 45.35
LLM-QAT 69.50 75.40 64.60 66.00 43.80 69.20 64.75
PEQA (our impl.) 72.97 77.80 67.72 67.13 40.27 73.35 66.54
LR-QAT (ours) 73.64 77.91 67.56 69.28 41.30 73.25 67.16
4-4-4 RTN 50.49 64.25 52.41 48.27 30.12 52.04 49.60
SmoothQuant 49.10 49.80 48.00 30.40 25.80 27.40 38.42
LLM-QAT 61.30 51.50 51.90 27.90 23.90 31.10 41.27
LLM-QAT + SQ 62.40 55.90 50.60 35.50 26.40 47.80 46.43
Outlier Suppression+ 60.21 62.73 52.96 39.98 30.29 44.39 48.43
OmniQuant§ 63.51 66.15 53.43 45.20 31.14 56.44 52.65
PEQA (our impl.) 65.69 72.31 59.83 56.52 34.22 61.79 58.39
LR-QAT (ours) 67.16 71.76 59.59 58.42 34.73 62.34 59.00
LLaMA-2 7B FP16 77.74 79.11 69.14 74.58 46.25 75.98 70.47
4-8-8 RTN 75.87 77.91 67.88 71.09 44.03 74.51 68.55
PEQA (our impl.) 77.37 77.97 69.77 70.54 43.52 75.50 69.11
LR-QAT (ours) 77.00 78.13 69.14 72.10 44.11 75.44 69.32
4-8-4 RTN 70.37 76.01 63.38 68.94 41.47 70.76 65.16
PEQA (our impl.) 74.71 77.48 67.40 69.28 42.75 73.75 67.56
LR-QAT (ours) 74.46 77.69 68.51 69.78 42.75 73.82 67.84
4-4-4 RTN 57.86 64.91 54.46 49.62 31.83 51.83 51.75
PEQA (our impl.) 67.09 70.67 60.06 54.80 32.17 62.76 57.93
LR-QAT (ours) 66.94 71.98 60.77 57.20 33.87 63.10 58.98
LLaMA-2 13B FP16 80.55 80.52 72.22 77.44 48.98 79.38 73.18
4-8-8 RTN 79.24 79.27 70.01 75.51 48.29 76.92 71.54
PEQA (our impl.) 79.02 80.20 71.19 76.60 48.72 79.22 72.49
LR-QAT (ours) 81.10 79.76 71.35 77.36 50.60 78.91 73.18
4-8-4 RTN 77.25 76.61 66.69 67.72 41.13 72.98 67.06
PEQA (our impl.) 78.01 79.22 69.30 75.59 48.21 77.78 71.35
LR-QAT (ours) 78.59 79.54 70.80 75.29 48.04 77.64 71.65
4-4-4 RTN 62.60 67.90 53.20 57.32 34.13 55.25 55.07
PEQA (our impl.) 68.72 74.21 62.35 63.85 36.01 68.44 62.26
LR-QAT (ours) 70.64 73.88 63.14 61.78 38.14 68.31 62.65
Table A7: A comparison between min-max and the best range setting used for round-to-nearest (RTN) initialization for LLaMA and Mistral models. We report WikiText-2 test set perplexity (lower is better) and average zero-shot accuracy (higher is better). §Uses asymmetric weight quantization.
# Bits Range WikiText-2 perplexity $\downarrow$ Avg. zero-shot accuracy $\uparrow$
estimator L1-7B L2-7B L2-13B L3-8B M-7B L1-7B L2-7B L2-13B L3-8B M-7B
FP16 5.68 5.47 4.88 6.14 5.25 69.68 70.47 73.18 74.22 75.69
W4 pc best est. $L^{4}$ $L^{3.5}$ $L^{3.5}$ $L^{3.5}$ $L^{4}$ $L^{4}$ $L^{3.5}$ $L^{3.5}$ $L^{3.5}$ $L^{4}$
best 6.33 6.14 5.21 7.53 5.91 68.51 68.88 71.73 72.19 73.44
min-max 6.85 7.14 5.40 10.53 6.33 66.23 66.41 72.19 67.44 71.84
W4 g128 best est. $L^{5}$ min-max min-max $L^{4}$ $L^{5}$ $L^{5}$ min-max min-max $L^{4}$ $L^{5}$
best 6.05 5.78 5.04 6.96 5.49 68.93 69.75 72.94 72.30 75.07
min-max 6.08 5.78 5.04 6.99 5.51 68.96 69.75 72.94 72.95 74.98
W3 pc best est. $L^{3.5}$ $L^{3.5}$ $L^{5}$ $L^{3.5}$ $L^{4}$ $L^{3.5}$ $L^{3.5}$ $L^{5}$ $L^{3.5}$ $L^{4}$
best 12.88 26.73 8.71 34.10 9.49 54.66 43.87 55.01 47.46 64.58
min-max 2.4e4 1.9e4 2.3e3 1.6e5 3.2e3 36.02 35.71 37.85 35.78 36.78
W3 g128 best est. $L^{5}$ $L^{4}$ $L^{5}$ $L^{5}$ $L^{5}$ $L^{5}$ $L^{4}$ $L^{5}$ $L^{5}$ $L^{5}$
best 7.95 7.61 6.20 15.11 6.77 63.50 63.20 67.60 57.74 69.35
min-max 8.10 8.22 6.14 29.38 7.22 62.69 64.07 66.81 54.54 68.35
W2 pc§ best est. $L^{2.4}$ $L^{3}$ $L^{3}$ $L^{3.5}$ $L^{3.5}$ $L^{2.4}$ $L^{3}$ $L^{3}$ $L^{3.5}$ $L^{3.5}$
best 4.9e3 5.2e3 5.2e3 6.4e4 6.8e3 37.92 36.52 36.27 36.80 36.59
min-max 1.1e5 2.5e4 4.9e4 1.4e6 7.5e4 39.07 36.53 39.01 37.69 38.13
W2 g128§ best est. $L^{4}$ $L^{3.5}$ $L^{5}$ $L^{4}$ $L^{5}$ $L^{4}$ $L^{3.5}$ $L^{5}$ $L^{4}$ $L^{5}$
best 708 2.5e3 115.6 1.4e4 369 39.74 37.94 41.12 36.97 41.30
min-max 3.6e3 5.9e3 341 2.8e5 3.4e3 37.06 36.78 40.30 37.18 37.16
[…]
grad. accumulation steps
[…]
@@ Appendix B Experimental details
Learning rate for $\bm{A}$, $\bm{B}$ (INT4 / INT3 / INT2) $\{10^{-5},10^{-4},10^{-3},10^{-2}\}$
Learning rate for $\mathbf{s}$ (INT4 / INT3) $\{0{{\textsuperscript{*}}},10^{-5}\}$
Learning rate for $\mathbf{s}$ (INT2) $\{10^{-5},10^{-4}\}$
[…]
Learning rate for $\bm{W}$ (full-model QAT only) $\{1,5,10,50,100\}\cdot 10^{-5}$
@@
Learning rate schedule for $\mathbf{s}$ linear (with warmup)
Learning rate schedule for $\bm{W}$ (full-model QAT only) linear (with warmup)
@@
Weight decay for $\mathbf{s}$ 00
Weight decay for $\bm{W}$ (full-model QAT only) $0.1$
Adam $\left(\beta_{1},\beta_{2}\right)$ $\left(0.9,0.95\right)$
@@
Table B1: Common hyperparameters used for experiments. *Is equivalent to freezing the quantization scale / zero offset obtained after initial range estimation ($\mathbf{s}=\mathbf{s_{0}}$, $\mathbf{z}=\mathbf{z_{0}}$).
@@
We experiment with both weight-only and weight-activation quantization. The default settings are INT4 / INT3 / INT2 per-channel (denoted ‘pc’) and group-wise weight quantization with a group size of 128 (denoted ‘g128’). We use symmetric quantization, except the INT2 case, where we use asymmetric quantization (1), for a fair comparison with related work. We quantize all linear layers, except the classification head. In weight-activation quantization, defaults are INT4 per-channel weight and per-token activation quantization [14]. Following OmniQuant [55], we quantize all inputs to matmuls with exception of the softmax output and additionally quantize the KV-cache as in LLM-QAT [44].
@@
We implement our method in PyTorch [53] and use training and evaluation pipelines from HuggingFace libraries [20, 38, 64]. For zero-shot evaluation, we use the LM Evaluation Harness framework [19]. Specifically, we use lm_eval v0.4.2 and report acc_norm for tasks where it’s available (PIQA, ARC-e, ARC-c, HellaSwag) and otherwise acc (BoolQ and Winogrande).
@@
We apply our methods to all linear layers in the attention blocks (both in self-attention and in the feed-forward network). We only train low-rank auxiliary matrices $\bm{A}$, $\bm{B}$ and the quantization parameters $\mathbf{s}$. Specifically, we freeze embedding layers, the final classification heads and RMSNorm parameters. In the case of asymmetric weight quantization, a zero offset $\mathbf{z}$ is set during range estimation phase and kept frozen throughout training.
We train on a small subset of SlimPajama [57], which is a close open-source replica of the dataset used for pre-training LLaMA models. We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set [^4] Specifically, we use the English subset of Wiki40b, https://huggingface.co/datasets/wiki40b, that contains cleaned-up text of English Wikipedia and training/validation splits. (512 sequences). For all weight-only and weight-activation quantization results, including the comparison with full-model QAT in Section 4.3, we train for $10^{4}$ steps. For ablation studies in Sections 4.1 and 4.2 we use shorter training of $10^{3}$ steps. Since the full-model QAT experiment requires more than 80GB of GPU memory, we apply checkpointing on the quantization function $\widehat{\bm{W}}$ to be able to run the experiment on an Nvidia A100 80GB GPU. We have also experimented with CPU optimizer state offloading, but that turned out to be significantly slower, see details in Table A8. The rest of the hyperparameters and their search spaces are listed in Table B1.
@@
All the experiments were executed on a single Nvidia A100 GPU equipped with 80GB of VRAM. Models of size 7B, 8B and 13B needed respectively around 2.25, 2.75 and 4.5 days for $10^{4}$ training steps experiments. For obtaining all the results in the paper, including the ablations, we needed 202 GPU days (A100). Including preliminary experiments that did not make it in the final paper and hyperparameter turning we estimate the total compute costs of this research to approximately 750 GPU days.